OpenStudy (anonymous):

how do you solve this problem its a proportion 2x/5=9/15

OpenStudy (anonymous):

cross multiply 2x*15=5*9 30x=45 x=45/30=3/2=1.5

OpenStudy (anonymous):

2x*5=5*3 10x=15 x=15/10=3/2=1.5 is that what you need?

OpenStudy (anonymous):

yeah i got the answer thanks